Find the potential difference across a resistor of resistance 10 ohm , if 10 to the power 20 electrons flow through it in one second.

Grade:10

3 Answers

Hrishikesh Sajan
6 years ago
v=IR
Divyadeep
6 years ago
charge on a single electron=1.6 x 10^-19C
so  charge on 10^20 electrons=[1.6 x 10^-19 x 10^20]C
= 16C
I(current)= net charge/time
A.T.Q....I = 16C/1second  = 16 A
so, current flowing through the resistor=16A
resistance value of the resistor  = 10 ohm
therefore by ohm’s law i.e. V=IR
Potential difference across the 10 ohm resistor= 16A x 10 ohm
=160 V
Akachdeep Medhi
6 years ago
Given,
No. of e- = 10^20
Charge of each e- = 1.6x10^(-19) C
So, charge across the resistor = 1.6x10^(-19)x10^20
= 1.6 x 10
= 16 C
Resistance = 10 ohm
So, by Ohm’?s Law
V = IR
=> V = 16 x 10
=> V = 160 V(Ans)